OpenStudy (anonymous):

Divide and simplify: 11v-55/21 that is divided by v-5/14v

OpenStudy (anonymous):

not sure how to type this, if know to maybe clarify?

OpenStudy (dumbcow):

\[\frac{\frac{11v-55}{21}}{\frac{v-5}{14v}}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[11v-55/21\div v-5/14v\]

OpenStudy (dumbcow):

factor the 11v-55 = 11(v-5) when dividing fractions, Flip and multiply \[=\frac{11(v-5)}{21}*\frac{14v}{v-5}\]

OpenStudy (dumbcow):

\[=\frac{22v}{3}\]

OpenStudy (anonymous):

wow, i'm glad someone's good at this stuff, Because I suck!!! thank u

OpenStudy (dumbcow):

np